find two consecutive even integers whose product is
288

Answer by kirubakaran(5)

find two consecutive even integers whose product is
288
let x and x+2 are the two even intergers
therefore, x (x+2) = 288 ( the product of two integers)
x^2 +2x =288
x^2+2x-288=0

x^2 + 18x -16x -288 =0
x (x+18)-16(x+18)=0
x =18 and x =16 are the the integers